OpenStudy (anonymous):

What is the solution set of |x - 6| = 4?

OpenStudy (anonymous):

(x-6) = 4 or -(x-6) = 4

OpenStudy (anonymous):

|x - 6| = 4?

OpenStudy (anonymous):

yeah the terms inside of the absolute bracket always has two solutions, positive and negative

OpenStudy (anonymous):

its a postive

OpenStudy (anonymous):

x-6 = 4 or -x+6 = 4 x-6+6 = 4+6 or -x+6-6 = 4-6 x = 10 or -x = -2 x = 10 or x = 2

OpenStudy (anonymous):

10-6 = 4, |4|=4 2-6=-4, |-4|=4 so that's the two solutions